6u+3v=-15
8u-5v=7
Solve system of equation by using substitution

Answer by Alan3354(69443)

6u+3v=-15 divide by 3 --> 2u + v = -5
8u-5v=7
Solve system of equation by using substitution
--------------
2u + v = -5 --> v = -2u - 5
8u - 5v = 7
------------
Sub for v in the 2nd equation
8u - 5*(-2u-5) = 7
Can you finish it?